Compute the impact on the money multiplier of an increase in the currency-to-deposit ratio from 10 percent to 15 percent when the reserve requirement is 10 percent of deposits, and banks’ desired excess reserves are 3 percent of deposits.

Instructions: Enter your responses rounded to two decimal places.

When desired currency holdings = 10% of deposits, m =

When desired currency holdings = 15% of deposits, m =
Answer #1

When Desired currency holdings = 10% of deposits , m= 1+1.1/ 1.1+0.1+0.03=1.71.

When Desired currency holdings = 15% of deposits , m = 1+1.15/1.15+0.1+0.03= 1.68.

Hope this helps.